In the following waveform, it takes 9 seconds for the wave to pass from point

A to point B.  


      A      B



What is the period of the wave? 

What is the frequency of the wave? 


2. Copy and complete the following table


PrefixAbbrevMultiplierMegaM103Centi10-3n

3. A 1.0 kg mass on a spring is stretched and released. The period of oscillation is measured to be 0.46 s. What is the spring constant?  

4. Describe how you would make a pendulum with a period of 1.0 sec. 


5. (a)What would be the period of a pendulum that is 1.0 m long? 

(b) Imagine you could travel to the moon where the acceleration due to gravity is 1.6 m/s2. What would be the period of a pendulum that is 1.0 m? 

6. (a) A power mower does 9.00 x 105 J of work in 0.500 h. What power does it develop?

(b) How long would it take a 500. W electric motor to do 1.50 x 105 J of work?



1
Expert's answer
2021-11-04T16:20:31-0400

The period is wavelength divided by speed:


"T=\lambda/v."

Frequency:


"f=\frac 1T."

3. The equation for period can help us to determine the spring constant:


"T=2\pi\sqrt{m/k},\\\space\\
k=\frac{4\pi^2 m}{T^2}=187\text{ N/m}."

4. Make a simple pendulum. Equation for period is


"T=2\pi\sqrt{\frac lg},\\\space\\
l=\frac{gT^2}{4\pi^2}=0.25\text{ m}."

So, take a string length of 25 cm.

5. (a) The period is


"T=2\pi\sqrt{\frac lg}=2.01\text{ s}."

(b) On the Moon:


"T=2\pi\sqrt{\frac lg}=4.97\text{ s}."

6. (a) Power:


"P=W/t=500\text{ W}."

(b) The time is


"t=\frac WP=300\text{ s}."